# The revised boomerang connectivity tables and their connection to the difference distribution table

### v2

@@ 7. Consequences of our results for some well-known functions
We now compute LBCT and UBCT entries of $F_{1}$. If $t=1$, then $F_{1}$ is APN, and using Corollary 6.1, we are done. Next, let us suppose that $t>1$ or equivalently, $\operatorname{DDT}_{F_{1}}(c,d)\neq 2$, then we will compute the LBCT and UBCT entries via (and using the notations of) Corollary 5.4. Obviously, $\operatorname{LBCT}_{F_{1}}(a,c,d)\neq 2$, for any $(a,c,d)\in\mathbb{F}_{2^{n}}^{*}\times\mathbb{F}_{2^{n}}^{*}\times\mathbb{F}_{2^{n}}^{*}$ and, $\operatorname{UBCT}_{F_{1}}(c,d,b)\neq 2$, for any $(c,d,b)\in\mathbb{F}_{2^{n}}^{*}\times\mathbb{F}_{2^{n}}^{*}\times\mathbb{F}_{2^{n}}^{*}.$ Notice that
$A_{k}=\left\{(c,c,d),(uc,c,d)\ \bigm{|}\operatorname{DDT}_{F_{1}}(c,d)=2^{t}=2k\mbox{~{}and~{} }u\in\mathbb{F}_{2^{t}}^{*}\setminus\{1\}\right\},$
as, in the case $\operatorname{DDT}_{F_{1}}(c,d)=2^{t}$, $(c,c,d)\in A_{k}$, for $b=d$ and $(uc,c,d)\in A_{k}$, for $b=(u+u^{2})c^{2^{s}+1}+ud$. Thus, from Corollary 5.4, we have $\operatorname{LBCT}_{F_{1}}(a,c,d)=2^{t}$, for $a\in\{c,uc\}$, where $u\in\mathbb{F}_{2^{t}}^{*}\setminus\{1\}$. This completes the LBCT case.
Similarly, one can write
$B_{k}=\left\{(c,d,d),(c,d,(u+u^{2})c^{2^{s}+1}+ud)\ |\operatorname{DDT}_{F_{1}}(c,d)=2^{t}\mbox{~{}and~{} }u\in\mathbb{F}_{2^{t}}^{*}\setminus\{1\}\right\},$
as, in the case $\operatorname{DDT}_{F_{1}}(c,d)=2^{t}$, $(c,d,d)\in B_{k}$, for $a=c$ and $(c,d,(u+u^{2})c^{2^{s}+1}+ud)\in B_{k}$, for $a=uc$. This gives us the UBCT of the Gold function. ∎
We will derive here the $\operatorname{FBCT}$ for the Gold function for any parameters (also found in [11], but our proof is significantly shorter).
@@
Let $F_{1}(X)=X^{2^{s}+1}$, with $\gcd(s,n)=t$ and $m=\frac{n}{t}$, where $ab\neq 0$. Without loss of generality, we take $b=1$. Then:
@@
This implies that, if $a\not\in\mathbb{F}_{2^{t}}^{*}$, then $\operatorname{FBCT}_{F_{1}}(a,1)=0$. Let us now assume that $a\in\mathbb{F}_{2^{t}}^{*}$. Note that the condition $\sum_{i=0}^{m-1}c^{2^{si}}=1$ is equivalent to $\operatorname{DDT}_{F_{1}}(1,c)=2^{t}$. Denoting as is usual $\omega_{i}=\left|\{c\bigm{|}\operatorname{DDT}_{F_{1}}(1,c)=i\}\right|$, $\operatorname{FBCT}_{F_{1}}(a,1)=\sum_{c\in\mathbb{F}_{2^{n}}}\operatorname{LBCT}_{F_{1}}(a,1,c)=2^{t}\omega_{2^{t}}=\sum_{i=0}^{2^{t}}i\omega_{i}=2^{n}$. This proves our claim. ∎
@@
Let $F$ be a differentially $4$-uniform function over $\mathbb{F}_{2^{n}}$ such that $\operatorname{DDT}_{F}(b,c)=4$ or $\operatorname{DDT}_{F}(b,c)=0$ for any $b,c\in\mathbb{F}_{2^{n}}$. For any $b,c$ such that $\operatorname{DDT}_{F}(b,c)=4$, denote by $\{y_{1},y_{1}+b,y_{2},y_{2}+b\}$ the four distinct solutions of the equation $F(X+b)+F(X)=c$. Then, if $y_{i}+y_{j}$, $i\neq j$, is independent of $c$, then by a similar argument, for $a\in\{y_{i}+y_{j},y_{i}+y_{j}+b\}$,
@@
Let $F_{2}(X)=X^{2^{2s}-2^{s}+1}$, where $\gcd(s,n)=2$, $n=2t$ and $t$ is odd and $3\nmid t$. Then for $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$, we have
$\operatorname{EBCT}_{F_{2}}(a,b,c,d)$ =$=\begin{cases}4&\mbox{ if }\operatorname{DDT}_{F_{2}}(c,d)=4\mbox{ and }a=c,b=d,\mbox{ or}\\ &\quad\operatorname{DDT}_{F_{2}}(c,d)=4,a=\omega c+\alpha^{2^{s}+1},\mbox{ and}b=\omega d+\alpha^{2^{3s}+1},\mbox{ or}\\ &\quad\operatorname{DDT}_{F_{2}}(c,d)=4,a=\omega^{2}c+\alpha^{2^{s}+1},\mbox{ and }b=\omega^{2}d+\alpha^{2^{3s}+1}\\ 0&\mbox{otherwise,}\end{cases}$
@@
$\operatorname{LBCT}_{F_{2}}(a,b,c)$ =$=\begin{cases}4&\mbox{ if }\operatorname{DDT}_{F_{2}}(b,c)=4\mbox{ and }a\in\left\{b,b\omega+\alpha^{2^{s}+1},b\omega^{2}+\alpha^{2^{s}+1}\right\},\\ 0&\mbox{otherwise,}\end{cases}$
@@
$\operatorname{UBCT}_{F_{2}}(a,b,c)$ =$=\begin{cases}4&\mbox{if}~{}\operatorname{DDT}_{F_{2}}(a,b)=4\mbox{ and }c\in\left\{b,b\omega+\alpha^{2^{3s}+1},b\omega^{2}+\alpha^{2^{3s}+1}\right\},\\ 0&\mbox{otherwise,}\end{cases}$
for $\alpha\in\mathbb{F}_{2^{n}}^{*}$ satisfying $a^{2^{2s}}+a^{2^{s}}\alpha^{2^{3s}-2^{s}}+a\alpha^{2^{3s}+2^{2s}-2^{s}-1}+b\alpha^{2^{2s}-1}=0$, ${\rm Tr}_{s}^{st}\left(1+\dfrac{a}{\alpha^{2^{s}+1}}\right)=0$ and $\omega$ is the cube root of unity.
@@
By Corollary 7.1, $\operatorname{EBCT}_{F_{2}}(a,b,c,d)$ equals
$\begin{cases}\operatorname{DDT}_{F_{2}}(c,d)&\mbox{ if }a=c,b=d,\mbox{ or}\\ &\quad\operatorname{DDT}_{F_{2}}(c,d)=4,a=z_{1}+z_{2},b=F_{2}(z_{1})+F_{2}(z_{2}),\mbox{ or}\\ &\quad\operatorname{DDT}_{F_{2}}(c,d)=4,a=z_{1}+z_{2}+c,b=F_{2}(z_{1})+F_{2}(z_{2}+c)\\ 0&\mbox{otherwise,}\end{cases}$
@@
Notice that the first equation of the above system has either 00 or $4$ solutions in $\mathbb{F}_{2^{n}}$, depending on whether $\sum_{i=0}^{t-1}\left(1+\dfrac{c}{\alpha^{2^{s}+1}}\right)^{2^{si}}$ is zero or not, respectively. This first equation is simply the equation for the DDT entry of $u^{2^{s}+1}$ at $(\alpha,c)$, and if $\operatorname{DDT}_{u^{2^{s}+1}}(\alpha,c)=4$, then the four solutions of $\left(\frac{u}{\alpha}\right)^{2^{s}}+\frac{u}{\alpha}+1+\frac{c}{\alpha^{2^{s}+1}}=0$ are $\{u,u+\alpha,u+\alpha\omega,u+\alpha\omega^{2}\}$. This implies that $u^{2^{s}+1},(u+\alpha)^{2^{s}+1},(u+\alpha\omega)^{2^{s}+1},(u+\alpha\omega^{2})^{2^{s}+1}$ are solutions of $F(X)+F(X+c)=d,$ or equivalently $\{z_{1},z_{1}+c,z_{2},z_{2}+c\}\subseteq\left\{u^{2^{s}+1},(u+\alpha)^{2^{s}+1},(u+\alpha\omega)^{2^{s}+1},(u+\alpha\omega^{2})^{2^{s}+1}\right\}$. Then for $z_{1}=u^{2^{s}+1}$, we have $z_{1}+c=(u+\alpha)^{2^{s}+1}$. Further, we split our analysis in the following two cases:
Case 1. Let $z_{2}=(u+\alpha\omega)^{2^{s}+1}$ Then $z_{1}+z_{2}=a$ and $F_{2}(z_{1})+F_{2}(z_{2})=b$ can be written as
$\begin{cases}u^{2^{s}+1}+(u+\alpha\omega)^{2^{s}+1}=a\\ u^{2^{3s}+1}+(u+\alpha\omega)^{2^{3s}+1}=b.\end{cases}$
@@
[20, Lemma 1] For any $X\in\mathbb{F}_{2^{n}}$ and $c\in\mathbb{F}_{2^{n}}$, let $t={\rm Tr}_{s}^{n}(c+1)\in\mathbb{F}_{2^{s}}$, where $n=4s$,

### v3

@@ 7. Consequences of our results for some well-known functions
We now compute LBCT and UBCT entries of $F_{1}$. If $t=1$, then $F_{1}$ is APN, and using Corollary 6.1, we are done. Next, let us suppose that $t>1$, then we will compute the LBCT and UBCT entries via (and using the notations and expressions of LBCT and UBCT) Theorem 5.1.
It is known from Theorem 5.1 that if $a=b,ab\neq 0$ or $a=0,b\neq 0$, then $\operatorname{LBCT}_{F_{1}}(a,b,c)=\operatorname{DDT}_{F_{1}}(b,c)=2^{t}$ if and only if ${\rm Tr}_{s}^{sm}\left(\frac{c}{b^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1)$; otherwise, it is zero. Now, let $a\neq b,ab\neq 0$ and $\operatorname{DDT}_{F_{1}}(b,c)=2k=2^{t}=4r$, where $r$ is a positive integer. Consider $S(b,c)=\{y_{1},y_{1}+b,y_{2},y_{2}+b,\ldots,y_{k},y_{k}+b\}$ to be the solution set of the equation $F_{1}(X+b)+F_{1}(X)=c$. Suppose that $y_{i_{1}}=X+v_{i_{1}}b$ and $y_{j_{1}}=X+v_{j_{1}}b$, where $X$ satisfies $X^{2^{s}+1}+(X+b)^{2^{s}+1}=c$, $v_{i_{1}}\neq v_{j_{1}}\in\mathbb{F}_{2^{t}}$ and $v_{i_{1}}+v_{j_{1}}\neq b$. Then, for $a=y_{i_{1}}+y_{j_{1}}=(v_{i_{1}}+v_{j_{1}})b$, System (3.2) has at least four solutions, namely, $\{(X+v_{i_{1}}b,X+v_{i_{1}}b+b),(X+v_{j_{1}}b,X+v_{j_{1}}b+b),(X+v_{i_{1}}b+b,X+v_{i_{1}}b),(X+v_{j_{1}}b+b,X+v_{j_{1}}b)\}$. We now need to investigate that whether System (3.2) has more than these four soultions, corresponding to $a=(v_{i_{1}}+v_{j_{1}})b$. From Theorem 5.1, it follows that $(y_{i_{2}},y_{i_{2}}+b)$, where $y_{i_{2}}=X+v_{i_{2}}b\in S(b,c)\setminus\{y_{i_{1}},y_{j_{1}},y_{i_{1}}+b,y_{j_{1}}+b\}$ and $v_{i_{2}}\in\mathbb{F}_{2^{t}}$, is a solution of System (3.2) if and only if there exists $y_{j_{2}}=X+v_{j_{2}}b\in S(b,c)\setminus\{y_{i_{1}},y_{j_{1}},y_{i_{1}}+b,y_{j_{1}}+b,y_{i_{2}},y_{i_{2}}+b\}$, $v_{j_{2}}\in\mathbb{F}_{2^{t}}$, such that $a=y_{i_{2}}+y_{j_{2}}$. Clearly, one can choose $y_{j_{2}}=X+(v_{i_{1}}+v_{i_{2}}+v_{j_{2}})$ to ensure $a=y_{i_{1}}+y_{j_{1}}=y_{i_{2}}+y_{j_{2}}$. Thus, we obtain another set of four solutions $\{(X+v_{i_{2}}b,X+v_{i_{2}}b+b),(X+v_{j_{2}}b,X+v_{j_{2}}b+b),(X+v_{i_{2}}b+b,X+v_{i_{2}}b),(X+v_{j_{2}}b+b,X+v_{j_{2}}b)\}$ for System (3.2) when $a=y_{i_{1}}+y_{j_{1}}$. Since $y_{i_{2}}\in S(b,c)\setminus\{y_{i_{1}},y_{j_{1}},y_{i_{1}}+b,y_{j_{1}}+b\}$ was choosen arbitrarily, it renders that every element in the set $S(b,c)$ is a solution of System (3.2) for $a=y_{i_{1}}+y_{j_{1}}$. Thus, $\operatorname{LBCT}_{F_{1}}(a,b,c)=\operatorname{DDT}_{F_{1}}(b,c)$.
We next compute the UBCT entries for $F_{1}$. It is evident from Theorem 5.1 that if $a=b=0$, then $\operatorname{UBCT}_{F_{1}}(a,b,c)=|F_{1}^{-1}(c+{\rm Im}(F_{1}))|$ where ${\rm Im}(F_{1})$ is the image of $F_{1}$ and $F_{1}^{-1}(\cdot)$ denotes the preimage of the argument. Moreover, if $b=c$ and $a\neq 0$ then $\operatorname{UBCT}_{F_{1}}(a,b,c)=\operatorname{DDT}_{F_{1}}(a,b)=2^{t}$ if and only if ${\rm Tr}_{s}^{sm}\left(\frac{b}{a^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1)$; otherwise, it is zero. Therefore, assume that $a\neq 0$ and $b\neq c$ and $\operatorname{DDT}_{F_{1}}(a,b)=2k=2^{t}=4r$, where $r$ is a positive integer. Define $S(a,b)=\{x_{1},x_{1}+a,x_{2},x_{2}+a,\ldots,x_{k},x_{k}+a\}$ as the solution set of the equation $F_{1}(X+a)+F_{1}(X)=b$. Suppose $x_{i_{1}}=X+w_{i_{1}}a$ and $x_{j_{1}}=X+w_{j_{1}}a$, where $w_{i_{1}}\neq w_{j_{1}}\in\mathbb{F}_{2^{t}}$, $w_{i_{1}}+w_{j_{1}}\neq a$ and $X^{2^{s}+1}+(X+a)^{2^{s}+1}=b$. Then, for $c=F_{1}(x_{i_{1}})+F_{1}(x_{j_{1}})=((w_{i_{1}}+w_{j_{1}})+(w_{i_{1}}+w_{j_{1}})^{2})a^{2^{s}+1}+(w_{i_{1}}+w_{j_{1}})b$, System (3.3) will have at least four solutions, namely, $\{(X+w_{i_{1}}a,X+w_{j_{1}}a),(X+w_{i_{1}}a+a,X+w_{j_{1}}a+a),(X+w_{j_{1}}a,X+w_{i_{1}}a),(X+w_{j_{1}}a+a,X+w_{i_{1}}a+a)\}$. Using the similar argument as in the case of EBCT, one can ensure that every element in the set $S(a,b)$ is a solution of System (3.3) for $c=F_{1}(x_{i_{1}})+F_{1}(x_{j_{1}})=((w_{i_{1}}+w_{j_{1}})+(w_{i_{1}}+w_{j_{1}})^{2})a^{2^{s}+1}+(w_{i_{1}}+w_{j_{1}})b$. Hence, we conclude that $\operatorname{UBCT}_{F_{1}}(a,b,c)=\operatorname{DDT}_{F_{1}}(a,b)$. ∎
We will derive here the $\operatorname{FBCT}$ for the Gold function for any parameters $a,b$ (also found in [12], but our proof is significantly shorter). As known, for monomials, one can take $b=1$.
@@
Let $F_{1}(X)=X^{2^{s}+1}$, with $\gcd(s,n)=t$ and $m=\frac{n}{t}$, where $a\in\mathbb{F}_{2^{n}}^{*}$. Then,
@@
This implies that, if $a\not\in\mathbb{F}_{2^{t}}^{*}$, then $\operatorname{FBCT}_{F_{1}}(a,1)=0$. Let us now assume that $a\in\mathbb{F}_{2^{t}}^{*}$. Note that the condition $\sum_{i=0}^{m-1}c^{2^{si}}=1$ is equivalent to $\operatorname{DDT}_{F_{1}}(1,c)=2^{t}$. Denoting as is usual $\omega_{i}=\left|\{c\bigm{|}\operatorname{DDT}_{F_{1}}(1,c)=i\}\right|$, $\operatorname{FBCT}_{F_{1}}(a,1)=\sum_{c\in\mathbb{F}_{2^{n}}}\operatorname{LBCT}_{F_{1}}(a,1,c)=2^{t}\omega_{2^{t}}=\sum_{i=0}^{2^{t}}i\omega_{i}=2^{n}$. This proves the claim. ∎
@@
Let $F$ be a differentially $4$-uniform function over $\mathbb{F}_{2^{n}}$ such that $\operatorname{DDT}_{F}(b,c)=4$ or $\operatorname{DDT}_{F}(b,c)=0$ for any $b,c\in\mathbb{F}_{2^{n}}$. For any $b,c$ such that $\operatorname{DDT}_{F}(b,c)=4$, denote by $\{y_{1},y_{1}+b,y_{2},y_{2}+b\}$ the four distinct solutions of the equation $F(X+b)+F(X)=c$. If $y_{i}+y_{j}$, $i\neq j$, is independent of $c$, then by using a similar argument as in the above corollary, for $a\in\{y_{i}+y_{j},y_{i}+y_{j}+b\}$, we obtain
@@
Let $F_{2}(X)=X^{2^{2s}-2^{s}+1}$ on $\mathbb{F}_{2^{n}}$, where $\gcd(s,n)=2$, $n=2t$ and $t$ is odd and $3\nmid t$. Then for $a,b,c,d\in\mathbb{F}_{2^{n}}$, we have
$\operatorname{EBCT}_{F_{2}}(a,b,c,d)$ =$=\begin{cases}2^{n}&\mbox{if }a=b=c=d=0,\\ 4&\mbox{if }\operatorname{DDT}_{F_{2}}(a,b)=4,a\neq 0\mbox{ and }c=d=0;\mbox{ or}\\ &~{}\operatorname{DDT}_{F_{2}}(c,d)=4,a=b=0\mbox{ and }c\neq 0;\mbox{ or}\\ &~{}\operatorname{DDT}_{F_{2}}(c,d)=4,ac\neq 0,a=c\mbox{ and }b=d;\mbox{ or}\\ &~{}\operatorname{DDT}_{F_{2}}(c,d)=4,a=\omega c+\alpha^{2^{s}+1},\mbox{ and }b=\omega d+\alpha^{2^{3s}+1};\mbox{ or}\\ &~{}\operatorname{DDT}_{F_{2}}(c,d)=4,a=\omega^{2}c+\alpha^{2^{s}+1},\mbox{ and }b=\omega^{2}d+\alpha^{2^{3s}+1},\\ 0&\mbox{otherwise,}\end{cases}$
@@
$\operatorname{LBCT}_{F_{2}}(a,b,c)$ =$=\begin{cases}2^{n}&\mbox{if }b=c=0,\\ 4&\mbox{if }\operatorname{DDT}_{F_{2}}(b,c)=4,a=0\mbox{ and }b\neq 0;\mbox{ or}\\ &~{}\operatorname{DDT}_{F_{2}}(b,c)=4,ab\neq 0\mbox{ and }a\in\left\{b,b\omega+\alpha^{2^{s}+1},b\omega^{2}+\alpha^{2^{s}+1}\right\},\\ 0&\mbox{otherwise,}\end{cases}$
@@
$\operatorname{UBCT}_{F_{2}}(a,b,c)$ =$=\begin{cases}2^{n}&\mbox{if }a=b=0,\\ 4&\mbox{if}~{}\operatorname{DDT}_{F_{2}}(a,b)=4,a\neq 0\mbox{ and }c\in\left\{b,b\omega+\alpha^{2^{3s}+1},b\omega^{2}+\alpha^{2^{3s}+1}\right\},\\ 0&\mbox{otherwise,}\end{cases}$
for $\alpha\in\mathbb{F}_{2^{n}}^{*}$ satisfying $a^{2^{2s}}+a^{2^{s}}\alpha^{2^{3s}-2^{s}}+a\alpha^{2^{3s}+2^{2s}-2^{s}-1}+b\alpha^{2^{2s}-1}=0$, ${\rm Tr}_{s}^{st}\left(1+\dfrac{a}{\alpha^{2^{s}+1}}\right)=0$ and $\omega$ is the primitive cube root of unity.
@@
Since $F_{2}$ is a permutation, Lemma  2.8 can be referred for determining EBCT, LBCT and UBCT entries of $F_{2}$ in cases where $abcd=0$. Thus, in what follows, we will assume that $abcd\neq 0$. By Corollary 7.1, for $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$, $\operatorname{EBCT}_{F_{2}}(a,b,c,d)$ equals
$\begin{cases}\operatorname{DDT}_{F_{2}}(c,d)&\mbox{if }a=c\mbox{ and }b=d;\mbox{ or}\\ &~{}\operatorname{DDT}_{F_{2}}(c,d)=4,a=z_{1}+z_{2}\mbox{ and }b=F_{2}(z_{1})+F_{2}(z_{2});\mbox{ or}\\ &~{}\operatorname{DDT}_{F_{2}}(c,d)=4,a=z_{1}+z_{2}+c\mbox{ and }b=F_{2}(z_{1})+F_{2}(z_{2}+c),\\ 0&\mbox{otherwise,}\end{cases}$
@@
Notice that the first equation of the above system has either 00 or $4$ solutions in $\mathbb{F}_{2^{n}}$, depending on whether $\sum_{i=0}^{t-1}\left(1+\dfrac{c}{\alpha^{2^{s}+1}}\right)^{2^{si}}$ is zero or not, respectively. This first equation is simply the equation for the DDT entry of $u^{2^{s}+1}$ at $(\alpha,c)$, and if $\operatorname{DDT}_{u^{2^{s}+1}}(\alpha,c)=4$, then the four solutions of $\left(\frac{u}{\alpha}\right)^{2^{s}}+\frac{u}{\alpha}+1+\frac{c}{\alpha^{2^{s}+1}}=0$ are $\{u,u+\alpha,u+\alpha\omega,u+\alpha\omega^{2}\}$. This implies that $u^{2^{s}+1},(u+\alpha)^{2^{s}+1},(u+\alpha\omega)^{2^{s}+1},(u+\alpha\omega^{2})^{2^{s}+1}$ are solutions of $F_{2}(X)+F_{2}(X+c)=d,$ or equivalently $\{z_{1},z_{1}+c,z_{2},z_{2}+c\}\subseteq\left\{u^{2^{s}+1},(u+\alpha)^{2^{s}+1},(u+\alpha\omega)^{2^{s}+1},(u+\alpha\omega^{2})^{2^{s}+1}\right\}$. Then for $z_{1}=u^{2^{s}+1}$, we have $z_{1}+c=(u+\alpha)^{2^{s}+1}$. Further, we split our analysis in the following two cases:
Case 1. Let $z_{2}=(u+\alpha\omega)^{2^{s}+1}$ Then $z_{1}+z_{2}=a$ and $F_{2}(z_{1})+F_{2}(z_{2})=b$ can be written as $u^{2^{s}+1}+(u+\alpha\omega)^{2^{s}+1}=a$ and $u^{2^{3s}+1}+(u+\alpha\omega)^{2^{3s}+1}=b$.
@@
[21, Lemma 1] For any $X\in\mathbb{F}_{2^{n}}$ and $c\in\mathbb{F}_{2^{n}}$, let $t={\rm Tr}_{s}^{n}(c+1)\in\mathbb{F}_{2^{s}}$, $n=4s$.